If a vulnerability exists in 480 servers out of the first 900 tested, what is the probability of NOT finding the vulnerability in the 901st server?
Answer #1

from given empriical data:

Probability of finding vulnerability in the 901st server =480/900 =8/15

therefore probability of NOT finding the vulnerability in the 901st server =1-8/15 =7/15